# Token-Regulated Group Relative Policy Optimization for Stable Reinforcement Learning in Large Language Models

### v1

@@ 3 Methodology
To realize what low/high-probability tokens are, in Figures 1, we visualize the top 100 low/high-probability tokens in each category. Figure 1(a) shows that high-probability tokens are often critical mathematical or logical symbols and formatting markers (e.g., operators, brackets, variable names). These elements are semantically indispensable: even a single error can invalidate an entire solution, meaning they are hardly replaceable. In contrast, Figure 1(b) illustrates that low-probability tokens are mostly generic, content words such as “output,” “particular,” or “location,” which contribute less to the core logical meaning of a sentence and can usually be substituted without altering the semantics. This contrast again provides intuition for why emphasizing high-probability tokens is more sensible—prioritizing them ensures stable updates on the tokens that carry the most critical signal.
We now investigate the derivative of the objective function in our approach, which has the following form
$\nabla_{\theta}J_{\text{TR-GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}}$ (12)
$\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\underbrace{\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\hat{A}_{i,t}\cdot\mathbb{I}_{\text{trust}}\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)+\beta\frac{\pi_{\text{ref}}(o_{i,t})}{\pi_{\theta}(o_{i,t})}-\beta\right)}_{\gamma_{i,t}}w_{i,t}\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\right],$
where we have defined
$\mathbb{I}_{\text{trust}}\!\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)=\begin{cases}0,&\text{if }\hat{A}_{i,t}>0\text{ and }\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}>w_{i,t}^{-1}\left(1+\epsilon_{h}\right),\\[6.0pt] 0,&\text{if }\hat{A}_{i,t}<0\text{ and }\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}<w_{i,t}^{-1}\left(1-\epsilon_{l}\right),\\[6.0pt] 1,&\text{otherwise}.\end{cases}$
Denote $g_{i,t}=\gamma_{i,t}w_{it}\nabla_{\theta}\log\pi_{\theta}(o_{i,t})$ as the token-based derivative in Eq. (12) relevant to $o_{i,t}$, we have the following theorem.
Theorem 3.2.
Under the same assumptions in Theorem 3.1, we can bound $\|g_{i,t}\|_{2}$ as
$\frac{w_{i,t}\left(1-\pi_{\theta}\left(o_{i,t}\right)\right)}{\sqrt{L}}\left|\gamma_{i,t}\right|\sum_{l=1}^{L}\left(a^{W}a_{l}^{G}\prod_{i=l}^{L}a_{i}^{J}\right)\leq\|g_{i,t}\|_{2}\leq\sqrt{2}w_{i,t}\left(1-\pi_{\theta}\left(o_{i,t}\right)\right)\left|\gamma_{i,t}\right|\sum_{l=1}^{L}\left[b^{W}b_{l}^{G}\prod_{i=l}^{L}b_{i}^{J}\right].$ (13)
According to the bounds in (13), for low-probability tokens, $1-\pi_{\theta}\left(o_{i,t}\right)$ is high and $w_{i,t}$ is low, whereas for high-probability tokens, $1-\pi_{\theta}\left(o_{i,t}\right)$ is low and $w_{i,t}$ is high. As a result, the product $w_{i,t}\left(1-\pi_{\theta}\left(o_{i,t}\right)\right)$ becomes more stable (i.e., exhibits lower variance) for both low-probability and high-probability tokens. Therefore, in our approach, $\|g_{i,t}\|_{2}$ may have lower variance for both low-probability and high-probability tokens, leading to a lower variance in the gradient norm compared to GRPO. To demonstrate this theoretical finding, we plot the gradient norms during training under three RLVR settings (described in detail in Section 4) in Figure 3. Compared to GRPO, TR-GRPO exhibits noticeably smaller fluctuations and fewer outlier spikes, resulting in a smoother and more stable optimization trajectory that aligns with our theoretical analysis.
Figure 3: Gradient norm trajectories during training under GRPO vs. TR-GRPO across three RLVR settings. TR-GRPO consistently exhibits lower variability and fewer spikes than GRPO, indicating stabilized updates in accordance with the bound in Eq. (13).
In addition to gradient norm analysis, we also track training rewards on Qwen2.5-3B-Instruct across logic, math, and agentic settings. As shown in Figure 4, TR-GRPO consistently yields higher and more stable reward trajectories than GRPO, further supporting the advantages of our proposed method in maintaining and encouraging the generation of high-probability tokens to form high-rewarded responses.
Figure 4: Training Reward trajectories during training under GRPO vs. TR-GRPO across three RLVR settings. The figure shows that TR-GRPO consistently achieves higher reward throughout training, indicating more effective learning and supporting our claim that probability-aware token weighting improves reasoning ability.
@@ 4.1 Experimental Setup
For the logic data, we adopt the K&K (Knights and Knaves) logic puzzles introduced in (Xie et al., 2024). Each puzzle describes a set of people who always lie or always tell the truth, and the task is to determine their identities given a set of statements. For training, we use the same dataset splits as in (Yang et al., 2025c), which include synthetic instances of increasing difficulty from 3 to 7 (measured by the number of people per puzzle). The reward function checks both the output format (requiring explicit <think> and <answer> tags) and the correctness of the final assignment. During evaluation, we follow the official harness and report accuracy across all difficulty levels, as well as breakdowns by puzzle size. This setting emphasizes step-by-step deductive reasoning and tests whether the model can align reasoning traces with verifiable logical consistency.
We next consider mathematical reasoning tasks, where correctness can be verified automatically against gold-standard answers. Following  (Yang et al., 2025c), we train on datasets that combine symbolic manipulation and arithmetic word problems, using a binary reward signal: $1$ if the final boxed answer matches the reference solution and 0 otherwise. The training corpus includes diverse math reasoning prompts designed to encourage structured derivations rather than direct guessing. For evaluation, we adopt multiple benchmarks that are standard in math-focused RLVR research: Olympiad Bench  Hendrycks et al. (2021), Minerva  (Lewkowycz et al., 2022), MATH-500  (He et al., 2024), AMC 2022-2023 and AIME 2024. For the first three benchmarks, evaluation is conducted using greedy decoding. For the last two benchmarks, consistent with standard practice, we generate 16 responses per question and report the mean accuracy across these samples (avg@16). Notably, since AIME 2024 is an extremely challenging dataset, we also report pass@16, which counts a problem as solved if at least one of the 16 responses is correct.
For agentic reasoning tasks, we examine an agentic setting that requires knowledge-intensive reasoning augmented with retrieval tools. Following the Search-R1  (Jin et al., 2025) configuration from VerlTool framework  (Jiang et al., 2025), the model interacts with an external tool server providing a retriever and other utilities such as Python or SQL. Training data consists of open-domain QA datasets where each instance requires grounding reasoning steps with retrieved evidence. The reward is based on exact-match correctness of the final answer, while intermediate tool calls are masked in the policy loss to avoid leakage of supervision. For evaluation, we follow the VerlTool benchmark and report exact-match accuracy on both General Q&A benchmarks (NQ  (Kwiatkowski et al., 2019), TriviaQA  (Joshi et al., 2017), PopQA (Mallen et al., 2022)) and multi-hop Q&A benchmark (HotpotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), MuSiQue (trivedi2022musique), Bamboogle (Press et al., 2022)).
@@ 4.2.2 Experiments on Math-related Datasets
Dataset Algorithms
Olympiad
Bench
Minerva
MATH
500
AMC
avg@16
AIME
pass@16
AIME
avg@16
Avg.
@@ Appendix A Related Work
Large language models (LLMs)  (Lambert et al., 2024; Gao et al., 2024; Team et al., 2025; Guo et al., 2025; Yang et al., 2025a) have recently achieved remarkable progress across a variety of natural language processing tasks. In particular, researchers are increasingly focused on enhancing their performance in reasoning-heavy domains such as mathematics (Cobbe et al., 2021; Hendrycks et al., 2024), coding (Jain et al., 2024), and scientific reasoning (Rein et al., 2024). rStar-Math (Guan et al., 2025) develops a self-evolved deep-thinking strategy that significantly boosts the reasoning capabilities of smaller LLMs. Snell et al. (Snell et al., 2024) introduce dense, process-based verifier reward models that adaptively refine a model’s response distribution based on the test-time prompt to strengthen reasoning skills. OpenAI’s O-series (Jaech et al., 2024) applies large-scale reinforcement learning to train models capable of solving complex reasoning tasks, achieving state-of-the-art results on multiple benchmarks.
@@
Before the emergence of reasoning-focused systems such as OpenAI’s O-series (Jaech et al., 2024), reinforcement learning (RL) had been widely applied in the form of reinforcement learning from human feedback (RLHF) to enhance large language models’ (LLMs) instruction-following and preference alignment (Ouyang et al., 2022). RLHF approaches are typically divided into online and offline optimization. Online algorithms—including PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), and REINFORCE (Williams, 1992)—update the model by generating outputs during training and receiving immediate reward signals. Offline variants such as DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), and KTO (Ethayarajh et al., 2024) optimize policies using pre-collected preferences from annotators or LLMs. While offline methods are generally more efficient, they often trail online training in terms of performance (Tang et al., 2024). More recently, reinforcement learning with verifiable rewards (RLVR) has been introduced as a promising paradigm for improving LLM reasoning, especially in domains like mathematics and programming. OpenAI o1 (Jaech et al., 2024) was the first to demonstrate that RL can successfully scale reasoning abilities. Subsequent models—including DeepSeek-R1 (Guo et al., 2025), Kimi-2 (Team et al., 2025), and Qwen3 (Yang et al., 2025a)—have either matched or surpassed its performance. Notably, DeepSeek-R1 highlights how strong reasoning can emerge from outcome-based optimization using online RL, particularly GRPO (Shao et al., 2024). These advances also inspired the “zero RL” paradigm, where reasoning capabilities are elicited directly from the base model without explicit RL fine-tuning. Building on this idea, follow-up work has proposed methods such as DAPO (Yu et al., 2025), VAPO (Yue et al., 2025), SimpleRLZoo (Zeng et al., 2025), and Open-Reasoner-Zero (Hu et al., 2025), which further explore RL-based reasoning at scale.
@@ B.1.1 Logic: Knights & Knaves (K&K)
Following  Xie et al. (2025); Yang et al. (2025c), we adopt LLMs after instruction fine-tuning (Qwen2.5-3B-Instruct  (Yang et al., 2025a) and Qwen2.5-7B-Instruct-1M  (Yang et al., 2025b)) as the initialization point. The tailored prompt designed for the LLMs is provided below.
@@
To promote chain-of-thought (CoT) reasoning in LLMs,  Xie et al. (2025) introduces a reward function with two main components, as shown in Table 5. The output is judged as fully correct if the LLMs generate CoT reasoning wrapped inside <think></think> tags, and the final prediction enclosed within <answer></answer> tags.
@@ B.1.2 Math-related dataset
As mentioned in Section 4.1, inspired by  (Yang et al., 2025c), we carry out additional experiments on two math-focused datasets, DSR-Uniform ($10,000$ problems, evenly covering difficulty levels) and ORZ ($57.000$ problems). In line with prior studies, we adopt Qwen2.5-7B  (Yang et al., 2025a) as the base model. In this setting, no instruction-tuned templates are applied; instead, we employ a simple prompt directly.
@@ B.1.3 Agentic: VT-Search (knowledge-augmented QA)
Question answering tasks frequently demand access to external knowledge that goes beyond a model’s parametric memory, especially for factual queries and multi-hop reasoning. To address this, we follow the Search-R1  (Jin et al., 2025) configuration from VerlTool framework  (Jiang et al., 2025) which incorporates a FAISS-based retrieval module, allowing agents to query a local knowledge base and extract the most relevant evidence for answering complex questions.
Building on prior work (Jin et al., 2025; Song et al., 2025), an E5 retriever (Wang et al., 2022) was employed with the 2018 Wikipedia dump (Karpukhin et al., 2020) as the indexed corpus. The agent alternates between retrieval operations and reasoning steps to form complete answers. we adopt Qwen2.5-3B  (Yang et al., 2025a) and Qwen3-4B-Instruct-2507  (Yang et al., 2025a) as the base models.
@@
For evaluation, we follow the VerlTool benchmark and report exact-match scores on both General Q&A benchmarks (NQ  (Kwiatkowski et al., 2019), TriviaQA  (Joshi et al., 2017), PopQA (Mallen et al., 2022)) and multi-hop Q&A benchmark (HotpotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), MuSiQue (trivedi2022musique), Bamboogle (Press et al., 2022))
@@ B.2 Hyperparameters
The key hyperparameter configurations for GRPO training are summarized in Table 6. We adopt the clip-higher technique from DAPO (Yu et al., 2025) to stabilize entropy and prevent collapse. Checkpoints are saved every 20 RL steps, and full implementation details are available in our code release. For token importance estimation (Eq. 11 in the main paper), we set $\alpha=2.0$ and $\mu=0.25$ with clipping bounds $L=1.0$ and $U=1.4$. The scaling parameter is fixed at $\tau=9.0$. All experiments are conducted on 8 × NVIDIA H100 GPUs.
@@ Appendix D Proof of Our Theories
Derivation of Derivative of GRPO in Eq. (3)
@@
Consider $A=\prod_{i=m}^{1}A_{i}$ and apply Inequality 15 recursively, we obtain
@@
Proof of Theorem 3.1
@@
Using the bound for $\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\|_{2}$ developed in Theorem 3.1, we reach the conclusion.

### v2

@@ 3.3 Token-Regulated GRPO for better generalization
Eq. (8) suggests that smaller gradient norms correspond to less sharp updates and hence better generalization. From Theorem 3.2, TR-GRPO suppresses extreme token gradients through the stabilizing factor $w_{i,t}(1-\pi_{\theta}(o_{i,t}))$. To validate this mechanism, we track gradient norms during training under three RLVR settings in Figure 3. Compared to GRPO, TR-GRPO exhibits smaller fluctuations and fewer outlier spikes, consistent with reduced sharpness and improved generalization.
Figure 3: Gradient norm trajectories during training under GRPO vs. TR-GRPO across three RLVR settings. TR-GRPO consistently exhibits lower variability and fewer spikes than GRPO, consistent with reduced sharpness in Eq. (8) and the bound in Eq. (14).
In addition to gradient-norm analysis, we also track training rewards across logic, math, and agentic settings. TR-GRPO yields higher and more stable reward trajectories than GRPO, indicating more effective learning while simultaneously controlling sharpness.
Figure 4: Training reward trajectories during training under GRPO vs. TR-GRPO across three RLVR settings. TR-GRPO achieves higher reward while also exhibiting lower sharpness as reflected by gradient norms.
@@ 4.1 Experimental Setup
For the logic data, we adopt the K&K (Knights and Knaves) logic puzzles introduced in (xie2024memorization). Each puzzle describes a set of people who always lie or always tell the truth, and the task is to determine their identities given a set of statements. For training, we use the same dataset splits as in (yang2025not), which include synthetic instances of increasing difficulty from 3 to 7 (measured by the number of people per puzzle). The reward function checks both the output format (requiring explicit <think> and <answer> tags) and the correctness of the final assignment. During evaluation, we follow the official harness and report accuracy across all difficulty levels, as well as breakdowns by puzzle size. This setting emphasizes step-by-step deductive reasoning and tests whether the model can align reasoning traces with verifiable logical consistency.
We next consider mathematical reasoning tasks, where correctness can be verified automatically against gold-standard answers. Following  (yang2025not), we train on datasets that combine symbolic manipulation and arithmetic word problems, using a binary reward signal: $1$ if the final boxed answer matches the reference solution and 0 otherwise. The training corpus includes diverse math reasoning prompts designed to encourage structured derivations rather than direct guessing. For evaluation, we adopt multiple benchmarks that are standard in math-focused RLVR research: Olympiad Bench  hendrycks2021measuring, Minerva  (lewkowycz2022solving), MATH-500  (he2024olympiadbench), AMC 2022-2023 and AIME 2024. For the first three benchmarks, evaluation is conducted using greedy decoding. For the last two benchmarks, consistent with standard practice, we generate 16 responses per question and report the mean accuracy across these samples (avg@16). Notably, since AIME 2024 is an extremely challenging dataset, we also report pass@16, which counts a problem as solved if at least one of the 16 responses is correct.
For agentic reasoning tasks, we examine an agentic setting that requires knowledge-intensive reasoning augmented with retrieval tools. Following the Search-R1  (jin2025search) configuration from VerlTool framework  (jiang2025verltool), the model interacts with an external tool server providing a retriever and other utilities such as Python or SQL. Training data consists of open-domain QA datasets where each instance requires grounding reasoning steps with retrieved evidence. The reward is based on exact-match correctness of the final answer, while intermediate tool calls are masked in the policy loss to avoid leakage of supervision. For evaluation, we follow the VerlTool benchmark and report exact-match accuracy on both General Q&A benchmarks (NQ  (kwiatkowski2019natural), TriviaQA  (joshi2017triviaqa), PopQA (mallen2022not)) and multi-hop Q&A benchmark (HotpotQA (yang2018hotpotqa), 2Wiki (ho2020constructing), MuSiQue (trivedi2022musique), Bamboogle (press2022measuring)).
@@ 4.2.2 Experiments on Math-related Datasets
Dataset Algorithms Olympiad Bench Minerva MATH 500 AMC avg@16 AIME pass@16 AIME avg@16 Avg.
@@ Appendix A Related Work
Large language models (LLMs)  (lambert2024tulu; gao2024designing; team2025kimi; guo2025deepseek; yang2025qwen3) have recently achieved remarkable progress across a variety of natural language processing tasks. In particular, researchers are increasingly focused on enhancing their performance in reasoning-heavy domains such as mathematics (cobbe2021training; hendrycks2024measuring), coding (jain2024livecodebench), and scientific reasoning (rein2024gpqa). rStar-Math (guan2025rstar) develops a self-evolved deep-thinking strategy that significantly boosts the reasoning capabilities of smaller LLMs. Snell et al. (snell2024scaling) introduce dense, process-based verifier reward models that adaptively refine a model’s response distribution based on the test-time prompt to strengthen reasoning skills. OpenAI’s O-series (jaech2024openai) applies large-scale reinforcement learning to train models capable of solving complex reasoning tasks, achieving state-of-the-art results on multiple benchmarks.
@@
Before the emergence of reasoning-focused systems such as OpenAI’s O-series (jaech2024openai), reinforcement learning (RL) had been widely applied in the form of reinforcement learning from human feedback (RLHF) to enhance large language models’ (LLMs) instruction-following and preference alignment (ouyang2022training). RLHF approaches are typically divided into online and offline optimization. Online algorithms—including PPO (schulman2017proximal), GRPO (shao2024deepseekmath), and REINFORCE (williams1992simple)—update the model by generating outputs during training and receiving immediate reward signals. Offline variants such as DPO (rafailov2023direct), SimPO (meng2024simpo), and KTO (ethayarajh2024kto) optimize policies using pre-collected preferences from annotators or LLMs. While offline methods are generally more efficient, they often trail online training in terms of performance (tang2024understanding). More recently, reinforcement learning with verifiable rewards (RLVR) has been introduced as a promising paradigm for improving LLM reasoning, especially in domains like mathematics and programming. OpenAI o1 (jaech2024openai) was the first to demonstrate that RL can successfully scale reasoning abilities. Subsequent models—including DeepSeek-R1 (guo2025deepseek), Kimi-2 (team2025kimi), and Qwen3 (yang2025qwen3)—have either matched or surpassed its performance. Notably, DeepSeek-R1 highlights how strong reasoning can emerge from outcome-based optimization using online RL, particularly GRPO (shao2024deepseekmath). These advances also inspired the “zero RL” paradigm, where reasoning capabilities are elicited directly from the base model without explicit RL fine-tuning. Building on this idea, follow-up work has proposed methods such as DAPO (yu2025dapo), VAPO (yue2025vapo), SimpleRLZoo (zeng2025simplerl), and Open-Reasoner-Zero (hu2025open), which further explore RL-based reasoning at scale.
@@ B.1.1 Logic: Knights & Knaves (K&K)
Following  xie2025logic; yang2025not, we adopt LLMs after instruction fine-tuning (Qwen2.5-3B-Instruct  (yang2025qwen3) and Qwen2.5-7B-Instruct-1M  (yang2025qwen2)) as the initialization point. The tailored prompt designed for the LLMs is provided below.
@@
To promote chain-of-thought (CoT) reasoning in LLMs,  xie2025logic introduces a reward function with two main components, as shown in Table 5. The output is judged as fully correct if the LLMs generate CoT reasoning wrapped inside <think></think> tags, and the final prediction enclosed within <answer></answer> tags.
@@ B.1.2 Math-related dataset
As mentioned in Section 4.1, inspired by  (yang2025not), we carry out additional experiments on two math-focused datasets, DSR-Uniform ($10,000$ problems, evenly covering difficulty levels) and ORZ ($57.000$ problems). In line with prior studies, we adopt Qwen2.5-7B  (yang2025qwen3) as the base model. In this setting, no instruction-tuned templates are applied; instead, we employ a simple prompt directly.
@@ B.1.3 Agentic: VT-Search (knowledge-augmented QA)
Question answering tasks frequently demand access to external knowledge that goes beyond a model’s parametric memory, especially for factual queries and multi-hop reasoning. To address this, we follow the Search-R1  (jin2025search) configuration from VerlTool framework  (jiang2025verltool) which incorporates a FAISS-based retrieval module, allowing agents to query a local knowledge base and extract the most relevant evidence for answering complex questions.
Building on prior work (jin2025search; song2025r1), an E5 retriever (wang2022text) was employed with the 2018 Wikipedia dump (karpukhin2020dense) as the indexed corpus. The agent alternates between retrieval operations and reasoning steps to form complete answers. we adopt Qwen2.5-3B  (yang2025qwen3) and Qwen3-4B-Instruct-2507  (yang2025qwen3) as the base models.
@@
For evaluation, we follow the VerlTool benchmark and report exact-match scores on both General Q&A benchmarks (NQ  (kwiatkowski2019natural), TriviaQA  (joshi2017triviaqa), PopQA (mallen2022not)) and multi-hop Q&A benchmark (HotpotQA (yang2018hotpotqa), 2Wiki (ho2020constructing), MuSiQue (trivedi2022musique), Bamboogle (press2022measuring))
@@ B.2 Hyperparameters
The key hyperparameter configurations for GRPO training are summarized in Table 6. We adopt the clip-higher technique from DAPO (yu2025dapo) to stabilize entropy and prevent collapse. Checkpoints are saved every 20 RL steps, and full implementation details are available in our code release. For token importance estimation (Eq. 12 in the main paper), we set $\alpha=2.0$ and $\mu=0.25$ with clipping bounds $L=1.0$ and $U=1.4$. The scaling parameter is fixed at $\tau=9.0$. All experiments are conducted on 8 × NVIDIA H100 GPUs.
@@ Appendix D Proof of Our Theories
Derivation of Derivative of GRPO in Eq. (3.2)
@@
Consider $A=\prod_{i=m}^{1}A_{i}$ and apply Inequality 16 recursively, we obtain
@@
Proof of Theorem LABEL:thm:grad_bound
@@
Using the bound for $\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\|_{2}$ developed in Theorem LABEL:thm:grad_bound, we reach the conclusion.